OpenStudy (luigi0210):

A particle travels along the curve y=sqr(x+4) When is the particle closest to the point (6,0)?

OpenStudy (anonymous):

when, or where?

OpenStudy (luigi0210):

When

OpenStudy (luigi0210):

Nobodies knows this?

OpenStudy (luigi0210):

anybody? :/ I really need help on this..

OpenStudy (anonymous):

OK, assuming I have the sense of the question then :we have some point (x0, y0) on the curve where the tangent to the curve at the point is orthogonal to the line from (6,0)?\[\frac{dy}{dx}=(1/2)(x+4)^{-1/2}\]\[\frac{dy}{dx}(x_{0})=\frac{1}{2\sqrt{x_{0}+4}}\]the product of the gradient of perpendicular lines is negative 1, so the slope from the point on the curve to (6,0) is\[-2\sqrt{x_{0}+4}\]equation of that line is: \[\frac{y_{0}-0}{x_{0}-6}= -2\sqrt{x_{0}+4}\]or knowing y0 from x0 :\[\frac{\sqrt{x_{0}+4}}{x_{0}-6}= -2\sqrt{x_{0}+4}\]which you can solve for x0.

OpenStudy (luigi0210):

Thank you very much! Would this also be used to find how far it is from (6,) at the moment?

OpenStudy (anonymous):

A Mathematica solution including a plot is attached.

OpenStudy (anonymous):

\[y=\sqrt{x+4}\]or\[y=(x+4)^{2}\]

OpenStudy (anonymous):

As for 'at the moment' then we'd need a time dependence for that.

OpenStudy (anonymous):

@hewsmike There is no sense of time expressed in the problem statement that I can see.

OpenStudy (anonymous):

@robtobey I'm not wanting to be rude but you've solved a different problem. :-)

OpenStudy (anonymous):

Second soluition based on\[y=\sqrt{(x+4)} \]